# Just Because You Can, Doesn’t Mean You Should: LLMs for Data Fitting

### v1

@@ 1 Introduction
Large language models (LLMs), as one of the most representative Generative AI applications, have been adopted across a wide variety of domains and tasks. While the majority of LLMs’ use cases are expectedly language-related (translation, Q&A, writing, ideation, etc.), there are also increasing uses of LLMs for tasks not apparently associated with language. We focus on one such practice, where LLM is used as an “off-the-shelf” data-fitting tool to generate predictions of interest. Data fitting refers to the task of using sets $\{(\boldsymbol{X},Y)\}$ of numerical input values $\boldsymbol{X}$ and target values $Y$ to learn (fit) a model $f(.)$ – i.e., $Y\sim f(\boldsymbol{X})$ – that can generate accurate predictions given inputs. For example, LLMs have been used for predicting property prices based on property features (e.g., number of bedrooms, bathrooms, etc.) and location information Tanlamai et al., (2024); performing time series forecasting and numeric prediction tasks in general by treating input data as a sequence of string tokens Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024); Vacareanu et al., (2024); detecting anomalies in time series data Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025); generating synthetic data in data-hungry domains Seedat et al., (2023); Wang et al., (2024); Brand et al., (2023); Isomura et al., (2024); and support causal effect estimation by predicting counterfactual outcomes (Huynh et al.,, 2025). In all of these examples, LLMs essentially act as a drop-in replacement for traditional, tabular supervised learning techniques to generate predictions based on input data (See SI Appendix Table S1 for more details about these examples).
@@
Our results reveal a fundamental challenge of using LLMs for data fitting: prediction robustness with respect to task-irrelevant variations, which is completely taken for granted when using tabular supervised learning techniques, can hardly be guaranteed in LLMs; at best, it can only be partially engineered in special-purpose foundation models. As an analogy, using LLMs for data fitting currently is like using a calculator for adding numbers, where the calculator produces different answers when given the same set of numbers in a different order. Therefore, we caution against the increasingly prevalent use of LLMs as a black-box data-fitting tool – the prediction quality cannot be ensured due to lack of robustness with respect to task-irrelevant variations.
@@ 2 Materials and Methods
We carry out comprehensive data-fitting experiments on GPT-4o-mini (a representative general-purpose close-weight LLM), [^3] Among many close-weight commercial LLMs, GPT-4o-mini is one of the most cost-effective flagship LLMs that support both completion and fine-tuning. Llama-3-8B-instruct (a representative general-purpose open-weight LLM), as well as TabPFN (currently best-performing tabular foundation model specifically trained for data-fitting tasks, Hollmann et al.,, 2025). We perform LLM-based data fitting via both in-context learning (ICL) and supervised fine-tuning (SFT). For GPT-4o-mini, we use OpenAI’s batch processing API for ICL and fine-tuning API for SFT. For Llama-3-8B-instruct, we download the model from Hugging Face [^4] Source: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. and carry out ICL using resources provided by Minnesota Supercomputing Institute. For the tabular foundation model (TabPFN), we follow TabPFN documentation to perform ICL using its Python implementation. The detailed prompts and procedures for both ICL and SFT are reported in SI Appendix, “LLM Prompts and Prediction Procedures” Section. Given the objective nature of the task, we use greedy decoding for all LLMs (i.e., selecting the next token as the one receiving highest predicted probability). This is done by setting LLM generation temperature to 0 and disabling random token sampling whenever possible. We also fix the random seed during SFT of LLMs as well as the training of all tabular supervised learning models to ensure reproducibility and fair comparisons across different models. All models are evaluated on the same 1,000 testing instances.
@@ 4.2 Prediction Sensitivity Cannot Be Avoided by Prompting
(iii) Row Order ++ Extra Inst. 1.2602 0.9498 0.8532
(44.03%) (23.09%) (8.96%)
The extra prompt instruction is clearly ineffective in countering prediction sensitivity induced by row order change. In configuration (iii), we still observe nontrivial sensitivity (up to 44.03%). Moreover, comparing configurations (ii) and (iii), we can see that inserting the extra instruction, which is intended to improve prediction robustness, can actually lead to higher prediction sensitivity in some settings (e.g., ICL500).
@@ 5.2 Attention-Based Mechanism Exploration
We observe similar “U-shaped” attention distributions under the ICL20 setting (with 20 ICL examples that are different from those in ICL10). This further indicates that the “U-shaped” pattern is associated with specific positions in the ICL prompt rather than the training examples we use. See SI Appendix, “Additional Analyses of Attention Distributions” Section for the detailed results.
The uneven attention distributions represent a plausible explanation for LLMs’ prediction sensitivity, because training examples and variables that happen to occupy the beginning / end positions in a prompt end up having larger influence in determining the output tokens. To illustrate this, we repeat the ICL10 experiment but shuffle the variable order. The “U-shaped” attention distributions persist, and variables that show up at beginning / end of each example by random chance now receive greater attentions from the output tokens. We include the results in SI Appendix, “Additional Analyses of Attention Distributions” Section.
@@ 6 Specialized Tabular Foundation Models Are Not Immune to Prediction Sensitivity
TabPNF achieves impressive predictive performance – with all 4,000 training examples, its average MAE clearly outperforms general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct) under both ICL and SFT. However, it is still not immune to task-irrelevant variations, with average prediction sensitivity ranging from 0.67% to 3.34%. Like the Llama model, re-running TabPFN with the same prompt yields exactly the same set of predictions, so the observed prediction sensitivity again cannot be attributed to token generation randomness.

### v2

# Robustness is Important: Limitations of LLMs for Data Fitting

@@ 1 Introduction
Large language models (LLMs), as one of the most representative Generative AI applications, have been adopted across a wide variety of domains and tasks. While the majority of LLMs’ use cases are expectedly language-related (translation, Q&A, writing, ideation, etc.), there are also increasing uses of LLMs for tasks not apparently associated with language. We focus on one such practice, where LLM is used as an “off-the-shelf” data-fitting tool to generate predictions of interest. Data fitting refers to the task of using sets $\{(\boldsymbol{X},Y)\}$ of numerical input values $\boldsymbol{X}$ and target values $Y$ to learn (fit) a model $f(.)$ – i.e., $Y\sim f(\boldsymbol{X})$ – that can generate accurate predictions given inputs. For example, LLMs have been used for predicting property prices based on property features (e.g., number of bedrooms, bathrooms, etc.) and location information Tanlamai et al., (2024); performing time series forecasting and numeric prediction tasks in general by treating input data as a sequence of string tokens Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024); Vacareanu et al., (2024); detecting anomalies in time series data Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025); generating synthetic data in data-hungry domains Seedat et al., (2023); Wang et al., (2024); Brand et al., (2023); Isomura et al., (2024); and support causal effect estimation by predicting counterfactual outcomes (Huynh et al.,, 2025). In all of these examples, LLMs essentially act as a drop-in replacement for traditional, tabular supervised learning techniques to generate predictions based on input data (See Appendix A for more details about these examples).
@@
Our results reveal a fundamental challenge of using LLMs for data fitting: prediction robustness with respect to task-irrelevant variations, which is completely taken for granted when using tabular supervised learning techniques, can hardly be guaranteed in LLMs; at best, it can only be partially engineered in special-purpose foundation models. As an analogy, using LLMs for data fitting currently can be viewed like using a calculator for adding numbers, where the calculator produces a different sum when given the same set of numbers in a different order. Therefore, we caution against the increasingly prevalent use of LLMs as a black-box data-fitting tool – the prediction quality cannot be ensured due to lack of robustness with respect to task-irrelevant variations.
@@ 2 Materials and Methods
We carry out comprehensive data-fitting experiments on GPT-4o-mini (a representative general-purpose close-weight LLM), [^3] Among many close-weight commercial LLMs, GPT-4o-mini is one of the most cost-effective flagship LLMs that support both completion and fine-tuning. Llama-3-8B-instruct (a representative general-purpose open-weight LLM), as well as TabPFN (currently best-performing tabular foundation model specifically trained for data-fitting tasks, Hollmann et al.,, 2025). We perform LLM-based data fitting via both in-context learning (ICL) and supervised fine-tuning (SFT). For GPT-4o-mini, we use OpenAI’s batch processing API for ICL and fine-tuning API for SFT. For Llama-3-8B-instruct, we download the model from Hugging Face [^4] Source: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. and carry out ICL using resources provided by Minnesota Supercomputing Institute. For the tabular foundation model (TabPFN), we follow TabPFN documentation to perform ICL using its Python implementation. The detailed prompts and procedures for both ICL and SFT are reported in Appendix B. Given the objective nature of the task, we use greedy decoding for all LLMs (i.e., selecting the next token as the one receiving highest predicted probability). This is done by setting LLM generation temperature to 0 and disabling random token sampling whenever possible. We also fix the random seed during SFT of LLMs as well as the training of all tabular supervised learning models to ensure reproducibility and fair comparisons across different models. All models are evaluated on the same 1,000 testing instances. Replication materials can be accessed at https://github.com/heziiiiiiiii/LLMs-for-data-fitting.
@@ 4.2 Prediction Sensitivity Cannot Be Avoided by Prompting
(iii) Row Order ++ Extra Inst. 1.2605 0.8900 0.8278
(44.06%) (15.33%) (5.71%)
The extra prompt instruction is clearly ineffective in countering prediction sensitivity induced by row order change. In configuration (iii), we still observe nontrivial sensitivity (up to 44.06%). Moreover, comparing configurations (ii) and (iii), we can see that inserting the extra instruction, which is intended to improve prediction robustness, can actually lead to higher prediction sensitivity in the ICL500 setting.
@@ 5.2 Attention-Based Mechanism Exploration
We observe similar “U-shaped” attention distributions under the ICL20 setting (with 20 ICL examples that are different from those in ICL10). This further indicates that the “U-shaped” pattern is associated with specific positions in the ICL prompt rather than the training examples we use. See Appendix C for the detailed results.
The uneven attention distributions represent a plausible explanation for LLMs’ prediction sensitivity, because training examples and variables that happen to occupy the beginning / end positions in a prompt end up having larger influence in determining the output tokens. To illustrate this, we repeat the ICL10 experiment but shuffle the variable order. The “U-shaped” attention distributions persist, and variables that show up at beginning / end of each example by random chance now receive greater attentions from the output tokens. We include the results in Appendix C.
@@ 6 Specialized Tabular Foundation Models Are Not Immune to Prediction Sensitivity
TabPNF achieves impressive predictive performance – with all 4,000 training examples, its average MAE clearly outperforms general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct) under both ICL and SFT. However, it is still not immune to task-irrelevant variations, with average prediction sensitivity ranging from 0.20% to 3.04%. Like the Llama model, re-running TabPFN with the same prompt yields exactly the same set of predictions, so the observed prediction sensitivity again cannot be attributed to token generation randomness.